# article.bib

### v2

@@ 2.5.4 Reward Function Design
To optimize the interaction between music rhythms and EEG signals for identifying and guiding individual perceptual states, this study designs a reward function based on user feedback and state transitions. The goal of the reward function is to encourage the system to guide subjects from their current state to the target state while penalizing state loss to prevent invalid data from affecting training.Reward Function Design Methods:
  • State-Matching Reward.The reward value is calculated based on the degree of match between the model’s predicted state and the subject’s confirmed state. The formula is:
@@
    where $p_{\text{model}}$ represents the state distribution predicted by the model (e.g., 70% summer, 30% spring), $p_{\text{user}}~$ represents the state distribution confirmed by the user (e.g., 80% summer, 20% spring), $\text{Loss}_{\text{transition}}$ describes the loss function for whether the state transition trend aligns with the user’s choice, and $\lambda 1,\lambda 2$ are parameters controlling the reward weights.
  • State Transition Reward.When the state transitions from "autumn" to "summer," the reward is proportional to the change in label distribution. The formula is:
    $R_{t}=\alpha\cdot\Delta S_{t-1\rightarrow t}-\beta\cdot 1(\text{im})$ (9)
    where $\Delta S_{t-1\rightarrow t}$ represents the absolute change in label distribution (e.g., from [80% autumn, 20% summer] to [60% autumn, 40% summer]), $1(\text{im})$ is an indicator for state loss (1 if state loss occurs, otherwise 0), and $\alpha$,$\beta$ are weights for reward and penalty.
  • Rhythm Adjustment Effect Reward.Reward is given based on changes in relaxation and focus scores after the subject clicks the "adapt to current state" button. The formula is:
    $R_{t}=\gamma\cdot(\Delta\text{Relaxation}+\Delta\text{Concentration})$ (10)
    where $\Delta\text{Relaxation}$ and $\Delta\text{Concentration}$ represent the changes in relaxation and focus scores, respectively. If the scores improve, the reward is positive; if they decrease, a negative reward is given.
@@ 2.5.6 Training Strategy
  • Exploration and Exploitation: Use an $\epsilon$-greedy strategy to balance exploration and exploitation. The initial $\epsilon$ value is 1.0 and gradually decays to 0.01 during training.
  • Parameter Updates: Use the Adam optimizer to update network parameters with a learning rate of 0.001.
  • Target Network Updates: Copy the current network parameters to the target network at regular intervals (e.g., every 10,000 steps).
@@ 3.1 CNN Model Perceptual Classification Performance Evaluation
The confusion matrix provides detailed information about the model’s classification performance across different categories. From the figure 4, it can be observed that the model performs well in distinguishing between "spring" and "summer" states, with only a small number of samples misclassified as "autumn". However, there is relatively more confusion between the "spring" and "winter" states, possibly due to the similarity in EEG features between these two seasons.
@@ 3.2 DQN Model Perceptual Guidance Performance Evaluation
On the validation set, the DQN model achieved a 92.45% success rate in guiding subjects’ EEG signals to the preset perceptual state. On average, it took 13.2 rhythm cycles (please replace with the actual average number of cycles) to complete a state guidance process, demonstrating the model’s good perceptual guidance capability and reasonable response speed.
@@
Although the EEG data shows that the DQN model achieved good results in perceptual guidance, subjective feedback from users revealed that approximately 50% of the researchers experienced psychological sensations corresponding to the target state during the rhythm adjustment process. This suggests that there is still room for improvement in the system’s effectiveness. Future enhancements could include introducing more personalized musical elements and more advanced reward functions.
@@ Conclusion
The above results demonstrate that the DQN model, through reinforcement learning, can effectively adjust music rhythms to guide EEG activity, thereby influencing users’ perceptual states to some extent [chandanwala2024hybridquantumdeeplearning]. On the validation set, the model achieved a high success rate in guiding perceptual states. Although there is still room for improvement in subjective experiences, the results validate the potential of this method in EEG biofeedback and perceptual guidance.
@@ 4.2 Evaluation of the Algorithm
The deep learning algorithm based on LOD theory proposed in this study demonstrates certain advantages in handling EEG analysis tasks for perceptual guidance. By introducing LOD theory into psychological modeling, we are able to approach the complex, multidimensional dynamic data of EEG signals from a new perspective. For example, in the model design, importance-based LOD technology enhances the model’s sensitivity to key EEG features by assigning higher weights to significant features, similar to focusing on important visual attributes in 3D modeling. The distance threshold method, applied to time-series pooling operations, simulates perceptual features at different time scales, allowing the density of EEG feature data to adjust dynamically with rhythm density, thereby better capturing the "outline" of perceptual scenes. Dynamic LOD technology, leveraging reinforcement learning, considers the dynamic nature of perceptual experiences in high-dimensional spaces and adjusts the level of detail based on real-time changes in the subject’s internal perspective, further optimizing model performance.
However, the algorithm also has some shortcomings. Current deep learning methods primarily focus on pattern matching at the sensory signal level. When dealing with psychological phenomena, due to a lack of deep understanding of the multi-level perceptual construction process, the predictive performance still has room for improvement. For instance, although we achieved some results in the experiments, there is still some confusion in distinguishing certain perceptual states (e.g., "autumn" and "winter"), possibly because we have not fully explored deeper perceptual information in EEG signals.
4.3 Psychological Phenomena Describable by the Algorithm
Perception, as the fundamental unit of psychological phenomena, has a unique multi-level structure. Sensation is the most basic neural signal, directly corresponding to data such as EEG, but it lacks higher-level psychological attributes. Perception, on the other hand, is the holistic interpretation of sensory signals, dependent on context and subjective experience [ma2024applying]. Emotions are composed of multiple perceptions, forming complex psychological states that can then be compressed into new perceptual units. Similarly, thinking is based on perception, and its results are also transformed into perceptual experiences. This multi-level relationship indicates that understanding the construction process of perception is crucial for accurately describing psychological phenomena.
4.4 Inspiration from Scaling Laws
In the era of large models, large language models rely on Scaling Laws (increasing data scale and model complexity). However, in the field of brain science, simply "stacking data and parameters" is not a viable strategy. The dimensionality of EEG data is not yet well-defined, and the intrinsic relationships within the data are holistic, making nonlinear patterns difficult to directly extract. This suggests that we need to find entry points to simplify complexity rather than blindly expanding data scale. LOD theory provides a perspective in this regard, capturing perceptual characteristics by "simplifying complexity" and extracting key patterns from massive data. This approach helps reduce reliance on large-scale annotated data while ensuring the robustness of perception, i.e., maintaining perceptual stability even when data is sparse or noisy [zhang2024investigatinggestaltprincipleclosure].
@@ 4.5.1 The Uncertainty Principle of Perception
Psychological phenomena (perception) are not solely determined by sensory input but are instead constructed through the interplay of the brain’s prior knowledge, subjective experience, and context. This characteristic is similar to the concept of superposition in quantum mechanics, where sensory data "collapses" into a specific psychological phenomenon only under a particular cognitive framework or perceptual perspective [rossi2024brain]. This reveals the subjectivity and unpredictability of human experience. Even if all neural signal data are available, without a "holistic framework," it is still impossible to reconstruct or predict individual psychological experiences, see Figure 1. Furthermore, our search for the causes of psychological phenomena may be misguided; psychological phenomena might be a priori existence, with neural activity serving as a post hoc rationalization.Although this perspective is somewhat paradigm-shifting, it provides a deeper direction for understanding psychological phenomena.
@@ 4.6.1 Future Improvements in Experimental Design
Although this study preliminarily validated the feasibility of guiding perceptual states by combining EEG biofeedback with music rhythm modulation, the experimental design still has some limitations. Specifically, in the selection of musical materials, the current drum rhythms, while effective in guiding users’ perceptual states, are limited by the simplicity of melodies and harmonies, which restricts the system’s expressive power. In future research, we plan to enrich the music library by introducing more diverse melodies and harmonic combinations to enhance the system’s guidance effectiveness [kim2024machine]. For example, by increasing the complexity of harmonies and the variability of melodies, the system can more precisely match users’ psychological states, thereby improving the user experience. Additionally, we aim to develop a more flexible user interface that allows users to customize music styles and rhythm parameters, further enhancing the system’s personalization and interactivity. We hope that these improvements will attract more researchers and users to explore the potential of EEG biofeedback and music modulation.

### v3

@@ 2.5.4 Reward Function Design
To optimize the interaction between music rhythms and EEG signals for identifying and guiding individual perceptual states, this study designs a reward function based on user feedback and state transitions. The goal of the reward function is to encourage the system to guide the subject from their current state to the target state while penalizing state loss to prevent invalid data from affecting training.
  • State-Matching Reward. The reward value is calculated based on the degree of match between the model’s predicted state distribution and the subject’s confirmed state distribution. The formula is:
@@
    where $p_{\text{model}}$ represents the state distribution predicted by the model (e.g., 70% summer, 30% spring), $p_{\text{user}}$ represents the state distribution confirmed by the user (e.g., 80% summer, 20% spring), $\text{Loss}_{\text{transition}}$ represents a penalty based on the alignment between the predicted state transition trend and the user’s desired direction. [Specify calculation, e.g., This is calculated as 0 if the transition aligns with the user’s implicit goal (moving towards target state) and 1 otherwise.], and $\lambda_{1},\lambda_{2}$ are parameters controlling the reward weights.
  • State Transition Reward. When the state transitions (e.g., from "autumn" to "summer"), the reward is proportional to the positive change in label distribution towards the target state, penalized by state loss. The formula is:
    $R_{t}=\alpha\cdot\Delta S_{t-1\rightarrow t}-\beta\cdot I_{\text{loss}}$ (9)
    where $\Delta S_{t-1\rightarrow t}$ represents the absolute change in label distribution towards the target state (e.g., increase in ’summer’ percentage when transitioning from autumn to summer), $I_{\text{loss}}$ is an indicator for state loss (1 if state loss occurs, otherwise 0), and $\alpha,\beta$ are weights for reward and penalty.
  • Rhythm Adjustment Effect Reward. Reward is given based on changes in Attention and Meditation scores after the subject clicks the "adapt to current state" button. The formula is:
    $R_{t}=\lambda_{3}\cdot(\Delta\text{Attention}_{t}+\Delta\text{Meditation}_{t})$ (10)
    where $\Delta\text{Attention}_{t}$ and $\Delta\text{Meditation}_{t}$ represent the changes in attention and meditation scores, respectively, following the user’s adaptation action. If the scores improve (increase), the reward is positive; if they decrease, a negative reward is given. $\lambda_{3}$ is a weighting parameter.
@@ 2.5.6 Training Strategy
  • Exploration and Exploitation: Use an $\epsilon$-greedy strategy to balance exploration (trying random actions) and exploitation (choosing the best-known action). The initial $\epsilon$ value is 1.0 (pure exploration) and gradually decays to 0.01 (mostly exploitation) over the course of training [Specify decay method, e.g., linearly over the first 100,000 steps].
  • Parameter Updates: Use the Adam optimizer to update network parameters based on the loss calculated from sampled experiences, with a learning rate of 0.001.
  • Target Network Updates: Copy the current network parameters ($\theta$) to the target network ($\theta^{-}$) at regular intervals (every 10,000 training steps) to stabilize learning.
@@ 3.1 CNN Model Perceptual Classification Performance Evaluation
The confusion matrix provides detailed information about the model’s classification performance across different categories. From the figure 4, it can be observed that the model performs well in distinguishing between "spring" and "summer” states, with only a small number of samples misclassified as "autumn". However, there is relatively more confusion between the "spring" and "winter" states, possibly due to the similarity in EEG features between these two seasons.
@@ 3.2 DQN Model Perceptual Guidance Performance Evaluation
On the validation set, the DQN model achieved a 92.45% success rate in guiding the subject’s EEG signals to the preset perceptual state. On average, it took 13.2 rhythm cycles to complete a state guidance process, demonstrating the model’s good perceptual guidance capability and reasonable response speed.
@@
Although the EEG data indicates successful state guidance by the DQN model, subjective feedback from the researcher revealed a more nuanced picture. While the system guided the EEG signals effectively towards the target patterns on the validation set, the researcher’s subjective feeling of being in the target perceptual state (e.g., "spring", "summer") varied considerably across different sessions and guidance attempts. This suggests a potential partial disconnect between the objective EEG markers targeted by the model and the emergent conscious experience of the individual. This observation highlights the complexity of linking neural activity to subjective states and underscores the need for further refinement to enhance the system’s effectiveness in reliably inducing the desired subjective experience. Future enhancements could include introducing more personalized musical elements reflecting the subject’s preferences and exploring more advanced reward functions that incorporate real-time subjective feedback or physiological indicators beyond basic EEG patterns.
@@ Conclusion
The above results demonstrate that the DQN model, through reinforcement learning, can effectively adjust music rhythms to guide EEG activity, thereby influencing the user’s perceptual states to some extent [chandanwala2024hybridquantumdeeplearning]. On the validation set, the model achieved a high success rate in guiding perceptual states. Although there is still room for improvement in subjective experiences, the results validate the potential of this method in EEG biofeedback and perceptual guidance.
@@ 4.1 Evaluation of the Algorithm
The deep learning algorithm based on LOD theory, as proposed in this study, demonstrated promising performance in EEG-based perceptual state analysis and guidance, achieving a classification accuracy of 94.05% and a guidance success rate of 92.45%. Several aspects of the LOD-inspired design likely contributed to these results. The importance-based LOD mechanism, by assigning higher weights to significant EEG features identified during training, may have enhanced the model’s sensitivity to discriminative patterns relevant to different perceptual states, thus boosting classification accuracy (1). Furthermore, the application of time-series pooling, analogous to the distance threshold method (2), potentially allowed the model to capture perceptual features at relevant temporal scales corresponding to the rhythm density, providing a robust representation. The dynamic LOD component (3), leveraging reinforcement learning and real-time context (Attention/Meditation scores), could adapt the feature representation dynamically, possibly contributing to the high success rate observed in the state guidance task by tailoring the process to the user’s immediate state. This integration of LOD principles appears advantageous for navigating the complexity of dynamic EEG signals compared to static feature extraction methods.
However, the algorithm also exhibits limitations, reflected in certain aspects of the results. The observed confusion between specific perceptual states (e.g., "autumn" and "winter" in the confusion matrix, Figure 4) suggests that the current feature extraction and LOD implementation may not fully capture the subtle nuances differentiating these subjective experiences from EEG signals alone. While the LOD framework aims to simplify complexity, the current realization might oversimplify or fail to extract deeper, more abstract perceptual information critical for finer distinctions. Additionally, the reported disconnect between objective guidance success and subjective experience (discussed in Results) further highlights this challenge. The algorithm, primarily focused on EEG pattern matching modulated by LOD, might not adequately model the multi-level construction process underlying subjective perception. This indicates that future iterations should focus not only on refining the LOD mechanisms but potentially incorporating richer contextual information or multimodal data to bridge the gap between neural signals and subjective reports.
4.2 Psychological Phenomena Addressed by the Algorithm
This study positions perception as the fundamental unit of analysis, moving beyond raw sensory signals like EEG oscillations. Psychological phenomena exhibit a multi-level structure: basic sensation (directly reflected in EEG features) lacks higher-level meaning, while perception represents the holistic interpretation of these signals, crucially dependent on context and subjective experience [ma2024applying]. Higher-level states like emotions and thoughts build upon this perceptual foundation. Our algorithm attempts to capture aspects of this perceptual level, rather than just classifying raw signal patterns.
The LOD-inspired approach, particularly mechanisms like importance-based feature weighting and dynamic context integration (using Attention/Meditation scores), aims to model this context-dependent, holistic nature of perception. For instance, the high classification accuracy achieved (94.05%) might suggest that the algorithm successfully identified EEG feature configurations corresponding to the distinct perceptual ’gestalts’ associated with the target states (spring, summer, autumn, winter). The dynamic LOD component, adjusting based on the user’s state, directly attempts to model the subjective and time-varying nature of the perceptual framework.
However, the difficulty in perfectly distinguishing certain states (e.g., autumn/winter) and the observed gap between objective EEG guidance and subjective experience highlight the challenges in algorithmically capturing the richness and subtlety of perception. Perception is not merely a complex pattern in EEG but an emergent phenomenon resulting from the interplay of sensory input, prior knowledge, and internal state [rossi2024brain]. Our current algorithm, while effective at pattern matching and dynamic adjustment based on limited context (rhythm, Attention/Meditation), likely only approximates the true perceptual construction process. Fully describing and predicting subjective perceptual experiences from EEG signals alone remains a significant hurdle, motivating future work to incorporate richer models of context and individual differences.
4.3 LOD as an Alternative to Brute-Force Scaling
The success of large language models often relies on Scaling Laws – increasing data scale and model complexity. Applying a similar "stacking data and parameters" approach directly to brain science, particularly EEG analysis, faces significant challenges. The dimensionality of EEG data in relation to psychological states is not well-defined, and the underlying relationships are often holistic and non-linear, making patterns difficult to extract through brute-force scaling alone, especially with limited datasets.
This study, utilizing data from a single subject (inherently limited in scale and diversity), exemplifies this challenge. Instead of pursuing massive data collection, we explored an alternative inspired by LOD theory: focusing on simplifying complexity and extracting key perceptual patterns. The LOD framework, by dynamically adjusting the level of detail processed based on context or feature importance, offers a strategy to potentially capture essential perceptual characteristics even from sparse or noisy data [zhang2024investigatinggestaltprincipleclosure]. This approach aims to reduce reliance on large-scale annotated datasets, which are often difficult and costly to obtain in psychological research, while attempting to maintain perceptual robustness. While the generalization capability of our model requires validation on larger datasets, the LOD-inspired methodology presents a potentially more data-efficient avenue for exploring complex neural signals related to subjective states compared to purely scale-driven approaches.
@@ 4.4.1 The Uncertainty Principle of Perception
Psychological phenomena (perception) are not solely determined by sensory input but are instead constructed through the interplay of the brain’s prior knowledge, subjective experience, and context. This characteristic is similar to the concept of superposition in quantum mechanics, where sensory data "collapses" into a specific psychological phenomenon only under a particular cognitive framework or perceptual perspective [rossi2024brain]. This reveals the subjectivity and unpredictability of human experience. Even if all neural signal data are available, without a "holistic framework," it is still impossible to reconstruct or predict individual psychological experiences, see Figure LABEL:fig:Figure06. Furthermore, our search for the causes of psychological phenomena may be misguided; psychological phenomena might be a priori existence, with neural activity serving as a post hoc rationalization.Although this perspective is somewhat paradigm-shifting, it provides a deeper direction for understanding psychological phenomena.
@@ 4.6.1 Future Improvements in Experimental Design
Although this study preliminarily validated the feasibility of guiding perceptual states by combining EEG biofeedback with music rhythm modulation, the experimental design still has some limitations. Specifically, in the selection of musical materials, the current drum rhythms, while effective in guiding the user’s perceptual states in this initial study, are limited by the simplicity of melodies and harmonies, which restricts the system’s expressive power. In future research, we plan to enrich the music library by introducing more diverse melodies and harmonic combinations to enhance the system’s guidance effectiveness, potentially allowing for more precise matching with the psychological states of diverse users [kim2024machine]. For example, by increasing the complexity of harmonies and the variability of melodies, the system could potentially more precisely match an individual user’s psychological state, thereby improving the user experience. Additionally, we aim to develop a more flexible user interface that allows users (in future extended studies) to customize music styles and rhythm parameters, further enhancing the system’s personalization and interactivity for a broader range of individuals. We hope that these improvements will attract more researchers and potential users to explore the potential of EEG biofeedback and music modulation.
@@ 4.6.2 Incorporating Melodies and Harmonies in Future Studies
Funding
This study was funded by the Science and Technology Plan Project of Inner Mongolia [Grant No. 2022YFSH0086] and the Inner Mongolia Medical University Joint Project [Grant No. YKD2023LH063]. We are grateful for their support.
Data Availability Statement
The raw dataset generated and analyzed during the current study is available in the Open Science Framework (OSF) repository at https://osf.io/qujz9/.
@@ 5 Abbreviation
RL Reinforcement Learning
ROC Receiver Operating Characteristic
SSE Screen Space Error